Please help me solve this equation using substitution
-3x-8y=20
-5x+y=19

Answer by Alan3354(69443)

Please help me solve this equation using substitution
-3x-8y=20
-5x+y=19
=============
Solve the 2nd eqn for y.
Sub for y in the 1st eqn.